DIVISION OF PARTNERSHIP EARNINGS

profit-and-loss ratio.

ATKINS, CAPITAL
Date Debit Credit Balance
Jan. 1 Balance . . . . . . . . . . . . . $162,000 $162,000
July 1 Additional Investment . . . 50,000 212,000
Sept. 3 Permanent reduction . . . . $30,000 $182,000
BRILL, CAPITAL
Date Debit Credit Balance
Jan. 1 Balance . . . . . . . . . . . . . $126,000 $126,000
COREY, CAPITAL
Date Debit Credit Balance
Jan. 1 Balance . . . . . . . . . . . . . $72,000 $72,000
Sept. 5 Additional investment. . . . 20,000 92,000

Net income is to be divided by allowing interest at 10% annually on the partners’ beginning of-
the-year investment capital and dividing the remainder 6/3/1 respectively.
a. The current year’s operations produced a $300,000 net income.
b. The current year’s operations produced an $80,000 net loss.

In this question I don't understand how they got the 6/3/1 ratio?

Can you please explain to me hoe to calculate this so I can see the 6/3/1 ratio?

Thanks, George. Alberta Canada
Answer #1

Answer

  • The profit sharing ratio is given in the question as 6:3:1. It is not to be calculated, it has already been provided.
  • However, on the basis of 6:3:1, the Net Income is to be distributed.
  • 6 + 3 + 1 = 10. Hence, Share of:
    --Atkings = 6 / 10
    --Brill = 3 / 10, and
    --Corey = 1 / 10
  • Check out below as to how 6/10, 3/10 and 1/10 is used to distribute the Net Income /(loss)
  • Requirement ‘a’

Net Income

Atkins

Brill

Corey

Distributed

To be distributed

Net Income

$300,000

Interest on Capital at 10%

$16,200

$12,600

$7,200

$36,000

$264,000

$264,000 shared in ratio of 6:3:1

$158,400 [264000 x 6/10]

$79,200 [264000 x 3/10]

$26,400 [264000 x 1/10]

$264,000

$0

Total Net Income distributed

$174,600

$91,800

$33,600

$300,000

  • Requirement ‘b’

Net Income

Atkins

Brill

Corey

Distributed

To be distributed

Net Income

($80,000)

0

$0

$0

$0

($80,000)

Interest on Capital at 10%

$16,200

$12,600

$7,200

$36,000

($116,000)

$(116,000) shared in ratio of 6:3:1

($69,600) =116000 x 6/10

($34,800) = 116000 x 3/10]

($11,600) = 116000 x 1/10

($116,000)

$0

Total Net Income distributed

($53,400)

($22,200)

($4,400)

($80,000)